OpenStudy (anonymous):

u=xy, v=x(y^4) solve for x and y in terms of u and v

OpenStudy (anonymous):

\[\large v = xy^4 \implies v = (xy) (y^3) \implies v = u(y^3)\] So, \[\large y^3 = \frac{v}{u} \implies y = \sqrt[3]{\frac{v}{u}}\] This is the value of y.. For x: \[\large u = xy \implies u^3 = x^3.y^3\] \[\large u^3 = x^3(\frac{v}{u}) \implies u^4 = v.x^3\] \[\large x^3 = \frac{u^4}{v} \implies x = \sqrt[3]{\frac{u^4}{v}} \implies x = u \sqrt[3]{\frac{u}{v}}\]